OpenStudy (anonymous):

Solve, using the substitution method. 5x + 2y = 7 y = x – 7 (Points : 1) (7, 0) (3, –4) (0, –7) (–1, –8)

HanAkoSolo (jamierox4ev3r):

since y is defined for us, we can plug it into the first equation

HanAkoSolo (jamierox4ev3r):

5x + 2 (x-7) =7 ^ see what i did there?

HanAkoSolo (jamierox4ev3r):

From here, distribute the two into the parentheses, and solve as a normal equation! Do you understand @crazy_girl121315 , or shall i explain more?

OpenStudy (mrnood):

the second equation gives you a way of writing y (that is x-7) you can 'substitiute that new way in the first equation so rewrite the first equation - wherever there is 'y' replace it with 'x-7'

HanAkoSolo (jamierox4ev3r):

5x + 2y = 7 ^ See that y there? we can replace it with the other equation, since that equation itself EQUALS y as @MrNood has said, we are "substituting a variable in the first equation, as it was defined in the second equation 5x + 2(x-7)=7

OpenStudy (anonymous):

im confuse

HanAkoSolo (jamierox4ev3r):

what don't you understand? Do you understand the substituting part? Or do you need help solving the actual equation?

OpenStudy (anonymous):

so the answer 0,7

OpenStudy (anonymous):

7,0

HanAkoSolo (jamierox4ev3r):

no that's not correct

OpenStudy (anonymous):

ok

HanAkoSolo (jamierox4ev3r):

Here, let me walk you through the process a little bit with this equation, 5x+2y=7 , it's impossible to solve for a variable since there are two variables present. So how do you get rid of the variable? Check if that variable can be replaced with another value. The first equation explicitly tells you that Y is equivalent to x-7. **x-7 will always equal y, even if we don't know what the exact value is at this point in time. Plugging in x-7 instead of y for this equation will give us like variables, and make us able to solve for x

HanAkoSolo (jamierox4ev3r):

So once you're here at 5x + 2(x-7) = 7, you distribute the two into the parentheses, like so:

HanAkoSolo (jamierox4ev3r):

distribution is basically just like multiplication

OpenStudy (anonymous):

thanks i got it

HanAkoSolo (jamierox4ev3r):

oh okay cool! What answer did you get?

OpenStudy (anonymous):

3, -4

HanAkoSolo (jamierox4ev3r):

yep that's right :)